OpenStudy (anonymous):

The number that completes this equation is 12. 2k2 + 6k + 180 = (k - 6)(k + __) True False need help please

OpenStudy (anonymous):

yes please im leaning towards false but wanna make sure

OpenStudy (anonymous):

ok working on it

OpenStudy (anonymous):

Thankyou love :)

OpenStudy (anonymous):

what lol

OpenStudy (anonymous):

yeah sorry thats what it is saying

OpenStudy (anonymous):

what grade are you in

OpenStudy (anonymous):

I just wanted to make sure thankyou so much for your help :)

OpenStudy (anonymous):

im new to this how do you fan you lol

OpenStudy (anonymous):

wow me to just click me

OpenStudy (anonymous):

Thankyou so much I greatly appreciate the help

sammixboo (sammixboo):

How to Fan Someone http://goo.gl/jL5V3G ANYWAYS, you first gotta plug in 12 in the blank, like so: 2k^2 + 6k + 180 = (k - 6)(k + 12) Now we gotta solve both sides of the equation.. We can first do this by simplifying the problem to this 2k^2 + 6k + 180 = k^2 + 6k - 72 Do you know what we do next?

OpenStudy (anonymous):

yes Sammi please

OpenStudy (anonymous):

if need help again chat with me ok

sammixboo (sammixboo):

Okie dokie, now we gotta subtract k^2 on both sides of the equation 2k^2 + 6k + 180 - k^2 = k^2 + 6k - 72 - k^2 k^2 + 6k + 180 = 6k - 72 Do you know what we do next?